OpenStudy (anonymous):

in need of assistance! Find the exact value for x if cos^2x=1/2sin 2x

OpenStudy (tkhunny):

\(\sin(2x) = 2\sin(x)\cos(x)\) -- That should help.

OpenStudy (anonymous):

Could you stay here so when I get the answer I can check it with you?

OpenStudy (anonymous):

X represents degrees, not a numerical answer, not sure if that effects it. . .

OpenStudy (tkhunny):

You should get over trying to make that a distinction for anything. It is entirely insignificant.

OpenStudy (anonymous):

Alrighty

OpenStudy (tkhunny):

Well, how's it going?

OpenStudy (anonymous):

I came up with x=0,90,45,225 degrees

OpenStudy (tkhunny):

Sure wish I knew how you came up with those. You didn't share much. If you had shared, I might be able to tell you how you managed 90º instead of the correct 180º.

OpenStudy (anonymous):

To be honest I'm not sure where I went wrong, I came up with cos x=o, but I was going through the steps my book displays.

OpenStudy (tkhunny):

\(\cos(x) = 0 \implies x = 90º\;or\;x = 270º\) You didn't fall for my little joke. 90º was good. 0º wasn't.

OpenStudy (anonymous):

Oh I see, I think I get it now, I just need help getting a problem started. I have two more similar problems would you be able to assist me with those ones? If you cannot I understand.

OpenStudy (tkhunny):

Post on a new thread. Tag me if you like.

OpenStudy (anonymous):

Alright, thank you! You're a life saver!

OpenStudy (tkhunny):

I've been called worse. :-)